OpenStudy (anonymous):

Solve the following system of equations graphically. y = 2x - 3 x + y = 3

OpenStudy (anonymous):

possible solutions are: {(2, 1)} {(1, 2)} {(-1, -2)} {(-2, -1)}

OpenStudy (campbell_st):

find the intercepts y = 2x - 3 x= 0 gives y-intercept y = -3 y = 0 gives x intercept x = 3/2 x + y = 3 x = 0 gives y-intercept y = 3 y = 0 gives x-intercept x = 3 plot (3/2, 0) and (0, -3) draw a line through the 2 points plot (0, 3) and (3, 0) draw a line through these 2 points... where the lines intersect is the solution to the 2 lines

OpenStudy (anonymous):

(2,1) ??

OpenStudy (campbell_st):

I haven't graphed the lines... I'm just giving you the method to graph them

OpenStudy (anonymous):

tha's it!! THank you!!!